OpenStudy (anonymous):

Solve the following systems by any method 3x + 2y = 5 4x + 5y = 16

OpenStudy (anonymous):

@AihberKhan This is my last one I promise :)

OpenStudy (shootingstar28):

Do I have to Graph it or solve it as a variable?

OpenStudy (anonymous):

@ShootingStar28 Solve as a variable :)

OpenStudy (anonymous):

Solve for x and y

OpenStudy (anonymous):

I'm not really sure, can you explain a little bit?

OpenStudy (anonymous):

3x / 3 = x and 2y + 5 / 3 = 6y + 5/3

OpenStudy (shootingstar28):

I will try to solve your system of equations. 3x+2y=5;4x+5y=16 Step: Solve3x+2y=5for x: 3x+2y+−2y=5+−2y(Add -2y to both sides) 3x=−2y+5 3x/ 3=−2y+5/3 (Divide both sides by 3) x=−2/3/y+5/3 Step: Substitute−2/3y+5/3 for x in 4x+5y=16: 4x+5y=16 4(−2/3y+5/3)+5y=167/3/y+20/3 =16(Simplify both sides of the equation) 7/3y+20/3+−20/3=16+−20/3(Add (-20)/3 to both sides) 7/3y=28/3/7/3y/7/3/=28/3/7/3(Divide both sides by 7/3) y=4 Step: Substitute 4 for y in x= −2/3y+5/3:x=−2/3y+5/3x=−2/3(4)+5/3x=−1(Simplify both sides of the equation) Answer: x=−1 and y=4

OpenStudy (shootingstar28):

You're welcome :)